# Harnessing The Collective Wisdom: Fusion Learning Using Decision Sequences From Diverse Sources

### v2

@@ 4.1 Connections to existing aggregation and derandomization procedures
    Ren and Barber (2024) show that the following is a set of generalized $e-$values.
@@
    We now explain how Equation (10) is related to Equation (1). Denote $\delta_{i}=\mathbb{I}(W_{i}\geq t_{\alpha})$. Then $t_{\alpha}$ can be written as
    $t_{\alpha}=\inf\left\{t>0:1+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-t)\leq\alpha\|\bm{\delta}\|_{0}\right\}.$
    Let $0<\breve{t}_{\alpha}<t_{\alpha}$ be such that $\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-\breve{t}_{\alpha})=1+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-t_{\alpha})$. We then have
@@
[…]
@@
    Li and Zhang (2023) show that the following is a set of generalized $e-$values.
@@
[…]
@@
4.2 Asymptotic FDR control
A key requirement for the validity of the IRT procedure is that the study-specific multiple testing procedure controls FDR at their pre-specified level $\alpha_{j}$. Theorems 2 and 3 implicitly assume that such an FDR control holds for finite samples, i.e. $\mathbb{E}[\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}/\|\bm{\delta}_{j}\|_{0}\vee 1]\leq\alpha_{j}$ for all $j\in[d]$. In reality, however, for some studies their FDR control may be asymptotic in $m_{j}$. In such a scenario, the IRT procedure in Algorithm 1 guarantees FDR control at level $\alpha$ as $m_{j}\to\infty$. We summarize the above discussion in the following proposition.
@@ 4.2 Asymptotic FDR control
Suppose study $j$ controls FDR at level $\alpha_{j}$ asymptotically, i.e., ${\rm FDR}(\bm{\delta}_{j})\leq\alpha_{j}+o_{p}(1)$. Then, Algorithm 1 ensures that FDR is controlled at level $\alpha$ asymptotically.
Proof of Proposition 1..
We first establish that $\mathbf{e}^{\tt agg}$ in Equation (3) is a set of generalized $e-$values asymptotically. Let $e_{i}^{\tt agg}$ be as defined in Equation (3). Akin to the proof of Theorem 2, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}^{\texttt{agg}}]$ $=\dfrac{m}{\sum_{j=1}^{d}m_{j}}\sum_{i\in\mathcal{H}_{0}}\sum_{j=1}^{d}\dfrac{m_{j}}{\alpha_{j}}\cdot\mathbb{E}\left[\dfrac{\delta_{ij}}{\|\bm{\delta}_{j}\|_{0}\vee 1}\right]\cdot\mathbb{I}(i\in\mathcal{M}_{j})$
$\leq\dfrac{m}{\sum_{i=1}^{m}m_{j}}\sum_{j=1}^{d}\frac{m_{j}}{\alpha_{j}}\cdot(\alpha_{j}+o_{p}(1))\leq m(1+o_{p}(1)).$
Next, we consider $e_{i}^{\tt agg*}$ from Equation (7). Similarly, following the same arguments in Lemma 1, we can show that $\mathbb{E}[e_{ij}]\leq\pi_{j}(1+o_{p}(1))$. Thus, using the same notation as in Theorem 3, we have $\mathbb{E}[\pi^{k}A_{i,\mathcal{S}_{ki}}]\leq 1+o_{p}(1)$. It follows that $\mathbf{e}^{\tt agg*}=\{e_{i}^{\tt agg*}\}_{i\in\mathcal{M}}$ is also a set of asymptotic generalized $e-$values. Let $\bm{\delta}=\{\delta_{1},\ldots,\delta_{m}\}$ be the decision of $e-$BH applied on a set of asymptotic generalized $e-$values $\mathbf{e}^{\tt agg}$. Note that $\delta_{i}=1$ indicates that $e_{i}^{\texttt{agg}}\geq m/(\alpha\|\bm{\delta}\|_{0}\vee 1)$ based on the decision rule of $e-$BH procedure. Thus, it holds that
@@
$\leq\frac{\alpha}{m}\cdot\mathbb{E}\left[\sum_{i=1}^{m}e_{i}^{\tt agg}\mathbb{I}(\theta_{i}=0)\right]=\frac{\alpha}{m}\cdot m(1+o_{p}(1))=\alpha+o_{p}(1),$
where the first inequality results from $\mathbb{I}(X>t)\leq X/t$ holds for all non-negative random variable $X$. ∎
4.3 When original $p-$values are available
While the IRT framework takes $\mathcal{D}_{j}$ as an input from each study $j$, it is important to consider the setting where study-specific $p-$values, denoted by $\{p_{ij}\}_{j\in\mathcal{N}_{i},i\in\mathcal{M}}$, are available. The goal under this setting is to aggregate the $p-$values pertaining to each hypothesis $i$ and then determine an appropriate threshold for FDR control at level $\alpha$ using the aggregated $p-$values. However, choosing an aggregation function for combining multiple $p-$values is challenging without making additional assumptions regarding their dependence structure (Vovk and Wang, 2020). Furthermore, if the underlying model is misspecified the validity of the corresponding $p-$values may be affected. In contrast, $e-$values are relatively more robust to such model misspecification (Wang and Ramdas, 2022) and particularly to dependence between the $p-$values (Vovk and Wang, 2021).
In this section, we consider the following calibrator from Equation (B.1) in Vovk and Wang (2021) for transforming $p_{ij}$ into corresponding $e-$values:
$e_{ij}^{\texttt{P2E}}(\kappa)=\begin{cases}\infty~{}&\text{if~{}}p_{ij}=0\\ \dfrac{\kappa(1+\kappa)^{\kappa}}{p_{ij}(-\log p_{ij})^{1+\kappa}}~{}&\text{if~{}}p_{ij}\in(0,e^{-\kappa-1}]\\ 0~{}~{}&\text{if~{}}p_{ij}\in(e^{-\kappa-1},1]\end{cases},$ (15)
where we choose $\kappa=1$ throughout the paper following the recommendation, and write $e_{ij}^{\texttt{P2E}}:=e_{ij}^{\texttt{P2E}}(1)$. Note that $e_{ij}^{\texttt{P2E}}$ as defined in Equation (15) are bonafide $e-$values. Therefore, to aggregate $e_{ij}^{\texttt{P2E}}$ we can simply take the average
${e}_{i}^{\texttt{{P2E},agg}}=\dfrac{1}{n_{i}}\sum_{j=1}^{d}e_{ij}^{\texttt{P2E}}~{}\mathbb{I}(i\in\mathcal{M}_{j}).$ (16)
We denote the method that applies the $e-$BH procedure on ${e}_{i}^{\texttt{{P2E},agg}}$ and $*{e}_{i}^{\texttt{{P2E},agg*}}$ as P2E and P2E, respectively. In Sections 5 and 6 we evaluate the numerical performance of P2E and P2E vis-a-vis IRT and IRT.

### v3

@@ D.1 Connections to existing aggregation and derandomization procedures
    Ren and Barber (2024) show that the following is a set of compound $e$-values.
@@
    We now explain how Equation (10) is related to Equation (3). Denote $\delta_{i}=\mathbb{I}(W_{i}\geq t_{\alpha})$. Then $t_{\alpha}$ can be written as $t_{\alpha}=\inf\left\{t>0:1+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-t)\leq\alpha\|\bm{\delta}\|_{0}\right\}.$ Let $0<\breve{t}_{\alpha}<t_{\alpha}$ be such that $\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-\breve{t}_{\alpha})=1+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-t_{\alpha})$. We then have
@@
    where the first inequality again follows from the definition of $t_{\alpha}$ in Equation (9). Hence, in the context of large-scale inference where $\|\bm{\delta}\|_{0}/(\|\bm{\delta}\|_{0}-\alpha^{-1})\overset{p}{\rightarrow}1$, Equation (3) and Equation (10) are asymptotically equivalent.
@@
    Li and Zhang (2023) show that the following is a set of compound $e$-values.
@@
    where the first inequality follows from Equation (11) and the second inequality follows from Equation (14). For large-scale inference problems, where $\|\bm{\delta}\|_{0}/(\|\bm{\delta}\|_{0}-\alpha^{-1})\overset{p}{\rightarrow}1$, Equation (3) and Equation (12) are asymptotically equivalent.
@@
D.2 Asymptotic FDR control
A key requirement for the validity of the IRT procedure is that the study-specific multiple testing procedure controls FDR at their pre-specified level $\alpha_{j}$. Theorems 2 and 3 implicitly assume that such an FDR control holds for finite samples, i.e. $\mathbb{E}[\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}/\|\bm{\delta}_{j}\|_{0}\vee 1]\leq\alpha_{j}$ for all $j\in[d]$. In reality, however, for some studies their FDR control may be asymptotic in $m_{j}$. In such a scenario, the IRT procedure guarantees FDR control at level $\alpha$ as $m_{j}\to\infty$. We summarize the above discussion in the following proposition.
@@ D.2 Asymptotic FDR control
Suppose study $j$ controls FDR at level $\alpha_{j}$ asymptotically, i.e., ${\rm FDR}(\bm{\delta}_{j})\leq\alpha_{j}+o_{p}(1)$. Then, IRT controls FDR at level $\alpha$ asymptotically.
Proof.
We first establish that $\mathbf{e}^{\tt agg}$ in Equation (4) is a set of compound $e$-values asymptotically. Let $e_{i}^{\tt agg}$ be as defined in Equation (4). Akin to the proof of Theorem 2, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}^{\texttt{agg}}]$ $=\dfrac{1}{d}\sum_{i\in\mathcal{H}_{0}}\sum_{j=1}^{d}\Bigl{\{}\dfrac{m_{j}}{\alpha_{j}}~\mathbb{E}\left[\dfrac{\delta_{ij}}{\max(\|\bm{\delta}_{j}\|_{0},1)}\right]~\mathbb{I}(i\in\mathcal{M}_{j})+\mathbb{I}(i\notin\mathcal{M}_{j})\Bigr{\}}$
$=\dfrac{1}{d}\sum_{j=1}^{d}\Bigl{\{}\dfrac{m_{j}}{\alpha_{j}}~{\rm FDR}(\bm{\delta}_{j})+\sum_{i\in\mathcal{H}_{0}}\mathbb{I}(i\notin\mathcal{M}_{j})\Bigr{\}}$
$\leq\dfrac{1}{d}\sum_{j=1}^{d}\Bigl{\{}m_{j}(1+o_{p}(1))+m-m_{j}\Bigr{\}}\leq m(1+o_{p}(1)).$
Next, we consider $e_{i}^{\tt agg*}$ from Equation (5). Similarly, following the same arguments in Lemma 1, we can show that $\mathbb{E}[e_{ij}]\leq(1/\pi_{j})(1+o_{p}(1))$. Thus, using the same notation as in Theorem 3, we have $\mathbb{E}[\pi^{k}A_{i,\mathcal{S}_{ki}}]\leq 1+o_{p}(1)$. It follows that $\mathbf{e}^{\tt agg*}=\{e_{i}^{\tt agg*}\}_{i\in\mathcal{M}}$ is also a set of asymptotic compound $e$-values. Let $\bm{\delta}=\{\delta_{1},\ldots,\delta_{m}\}$ be the decision of e-BH applied on a set of asymptotic compound $e$-values $\mathbf{e}^{\tt agg}$. Note that $\delta_{i}=1$ indicates that $e_{i}^{\texttt{agg}}\geq m/\{\alpha\max(\|\bm{\delta}\|_{0},1)\}$ based on the decision rule of e-BH procedure. Thus, it holds that
@@
$\leq\frac{\alpha}{m}\cdot\mathbb{E}\left[\sum_{i=1}^{m}e_{i}^{\tt agg}\mathbb{I}(\theta_{i}=0)\right]=\frac{\alpha}{m}\cdot m(1+o_{p}(1))=\alpha+o_{p}(1).$
Appendix E IRT for alternative forms of Type I error control
E.1 IRT for $k$-Family-wise error rate ($k$-FWER) control
$\mathbb{P}\Big{(}\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}\geq k_{j}\Big{)}\leq\alpha_{j},\quad\forall j\in[d].$ (15)
If study $j$ satisfies Equation (15) then the evidence indices in Equation (16) are compound $e$-values. To see this, let $V_{j}$ denote the number of false rejections made by study $j$. Then,
We note that if additional information, such as what procedure study $j$ used to control $k_{j}$-FWER, is available then it becomes feasible to devise more powerful compound $e$-values. For example, if Bonferroni procedure is used (i.e. $H_{0i}$ is rejected by study $j$ if and only if its $p$-value is $\leq k_{j}\alpha_{j}/m_{j}$), then we can verify that $e_{ij}={m_{j}\delta_{ij}}/{\alpha_{j}k_{j}}$ for all $i\in\mathcal{M}_{j}$, is also a compound $e$-value. To see this, observe that $\mathbb{E}[\sum_{i\in\mathcal{H}_{0j}}e_{ij}]\leq{m_{j}}(k_{j}\alpha_{j})^{-1}\mathbb{E}[\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}]={m_{j}}{(k_{j}\alpha_{j})^{-1}}\sum_{i\in\mathcal{H}_{0j}}\mathbb{P}(\delta_{ij}=1)\leq{m_{j}}({k_{j}\alpha_{j}})^{-1}\sum_{i\in\mathcal{H}_{0j}}{k_{j}\alpha_{j}}/{m_{j}}\leq m_{j}.$
Evidence aggregation. Since the evidence indices in Equation (16) are compound $e$-values, Equation (4) provides the evidence aggregation scheme in this setting and Theorem 2 guarantees that these aggregated evidences continue to be compound $e$-values associated with $\mathcal{H}_{0}$. Furthermore, if the conditions of Theorem 3 hold then Equation (5) represents the aggregated evidence indices.
$k$-FWER control. Given the compound $e$-values $\mathbf{e}^{\tt agg}=\{e_{1}^{\tt agg},\ldots,e_{m}^{\tt agg}\}$ from Step 2 above, IRT rejects $H_{0i}$ if and only if $e_{i}^{\tt agg}\geq m/(\alpha k)$. This procedure controls the $k$-FWER at level $\alpha$ since
E.2 IRT for Per-family error rate (PFER) control
Evidence construction. We consider the evidence index
$e_{ij}=\dfrac{m_{j}\delta_{ij}}{k_{j}},\quad\forall i\in\mathcal{M}_{j}.$ (17)
It is then straightforward to check that $\mathbf{e}_{j}=\{e_{ij}\}_{i\in\mathcal{M}_{j}}$ are a set of compound $e$-value associated with $\mathcal{H}_{0j}$.
Evidence aggregation. The evidence indices in Equation (17) are compound $e$-values. Therefore, Equation (4) continues to provide the evidence aggregation scheme in this setting and Theorem 2 guarantees that these aggregated evidences are compound $e$-values associated with $\mathcal{H}_{0}$. Furthermore, if the conditions of Theorem 3 hold then Equation (5) represents the aggregated evidence indices.
$k$-PFER control. Given the Genralized $e$-values $\mathbf{e}^{\tt agg}=\{e_{1}^{\tt agg},\ldots,e_{m}^{\tt agg}\}$ from Step 2 above, IRT rejects $H_{0i}$ if and only if $e_{i}^{\tt agg}\geq m/k$. This procedure controls the PFER at level $k$ since $\mathbb{E}[\sum_{i\in\mathcal{H}_{0}}\mathbb{I}\Big{(}e_{i}\geq m/k\Big{)}]\leq\mathbb{E}[\sum_{i\in\mathcal{H}_{0}}\mathbb{I}\Big{(}e_{i}k/m\geq 1\Big{)}]\leq\mathbb{E}[\sum_{i\in\mathcal{H}_{0}}(e_{i}k/m)]\leq{k}{m^{-1}}\mathbb{E}[\sum_{i\in\mathcal{H}_{0}}e_{i}]\leq k$.